A bag contains 8 red and 5 white balls. Three balls are selected successively with replacement and without replacement. Find the probability that:
i. Three white balls are selected with replacement. (5 Marks)
ii. Three red balls are selected without replacement. (5 Marks)

Answer by htmentor(1343)

There are a total of 13 balls in the bag.
i. With replacement, with each draw there is the same probability of getting a white ball.
Since there are 5 white balls P(white) = 5/13.
Thus P(3 white) = (5/13)(5/13)(5/13) = 0.057
ii. Without replacement, there are one fewer red and one fewer total balls with each draw
Thus P(3 red) = (8/13)(7/12)(6/11) = 0.196